OpenStudy (anonymous):

what is the value of y in the equation 4y-2(1-y)=-44

OpenStudy (anonymous):

1 Simplify 4y−2(1−y) to 2(3y−1) 2(3y−1)=−44 2 Divide both sides by 2 3y−1=−442 3 Simplify 442 to 22 3y−1=−22 4 Add 1 to both sides 3y=−22+1 5 Simplify −22+1 to −21 3y=−21 6 Divide both sides by 3 y=−213 7 Simplify 213 to 7 y=−7

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can i get a medal?

OpenStudy (anonymous):

i dont know how to give medals .. can you tell me & i'll give it to you . @ugh.math

OpenStudy (anonymous):

ugh idk either lol